loge (x-1) e 1 - logs &
Answer #1

lay, (2-1) = 1 - loga(z) log.cx-1) +logoca) 1 logo loge (x-1). - 2.(2-1) = 6 2 & -a X- 60 x² 3a+22-6=0 712-3) + 2(x-3)=0 (x+2